OpenStudy (anonymous):

VS=90V, R=1kΩ, and the transistor parameter f=1mAV3. The amplifier is biased with VIN=3V. What is the dc operating point (bias) output voltage VOUT (in Volts) of this amplifier? What is the small signal gain (vOUTvIN) of this amplifier when it is biased as above?

OpenStudy (anonymous):

63 -27

OpenStudy (anonymous):

Vout=Vs-(f*(VIN^3))*R Gain=-3*f*(VIN^2)*R